If 90 out of 200 individuals in a population express the recessive phenotype, what proportion of the population are expected to be heterozygotes?
Answer #1

1. frequency of recessive alleles = q^2 = 90/200 = 0.45

q = 0.671

frequency of recessive alleles = 0.671

2. frequency of recessive alleles = q ^ 2 = 90/200 = 0.45

q =0. 671

p + q = 1

p + 0.671 = 1

p = 0.329

2 p q = 2 (0.671* 0.329) = 0.44

expeted  heterozygotes population is 0.44

to check all are correct or not = p( square) + q ( square) + 2pq = 1

(p square = 0.329 square = 0.1082)

0.1082 + 0.451 + 0.441 = 1 ( approx)

2.